xx CapetIOScalar

CapetIOScalar

Erik Schnetter <schnetter@aei.mpg.de>

June 11, 2004

Abstract

1 Introduction

2 Utilities These scripts may be useful handling CarpetIOASCII data:

Carpet2ygraph.pl A perl script to convert to a format suitable for the xgraph or ygraph packages of P. Walker and D. Pollney.

mergeCarpetIOScalar.pl A perl script to remove duplicate datasets from file(s).

The first script resides in CarpetIOASCII/src/utils and was meant primarily to convert 1D data, since the CarpetIOScalar data format is easily readable also by ygraph. Anyway, this script converts also scalar data files to the original 2-column ygraph format. Usage: Carpet2ygraphCat.pl Inputfile.

The second script can be used to merge CarpetIOScalar output written before and after recovery. It reads one or more files in CarpetIOScalar format and writes their contents to STDOUT, eliminating duplicate datasets (all but the last occurance are discarded), which may be created when a run died abruptly without checkpointing at the last output itaration and is recovered from a previous checkpoint. Usage: mergeCarpetIOScalar.pl list

of files.
Example 1: mergeCarpetIOScalar.pl alp.norm1.asc > alp.norm1.asc.merged.
Example 2: mergeCarpetIOScalar.pl alp.norm1.asc.firstCHKPT alp.norm1.asc.secondCHKPT >
alp.norm1.asc.merged.

This script was written by Thomas Radke.

References

3 Parameters

11allxreductionsxinxonexfile 10Scope: privatex01  BOOLEAN

11Description: Write all requested reductions in one file instead of per reduction

11                  10  x01  Default: no

11

11onexfilexperxgroup 10Scope: privatex01  BOOLEAN

11Description: Write one file per group instead of per variable

11                  10  x01  Default: no

11

11outxprecision  10Scope: privatex01  INT

11Description: How many digits to output floating-point numbers with

11 Range 10 x01 Default: 15
110:*  10Number of precision digits

11outscalarxcriterion 10Scope: privatex01  KEYWORD

11Description: Criterion to select scalar output intervals, overrides outxevery

11 Range 10 x01 Default: iteration
11never  10Never output
11iteration  10Output every so many iterations
11divisor  10Output if iteration mod divisor == 0.
11time  10Output every that much coordinate time

11

11outscalarxdir  10Scope: privatex01  STRING

11Description: Name of scalar output directory, overrides outxdir

11 Range 10 x01 Default: (none)
11ˆ$  10Empty: use IO::outxdir
11.+  10Not empty: directory name

11

11outscalarxdt  10Scope: privatex01  REAL

11Description: How often to do scalar output, overrides outxdt

11 Range 10 x01 Default: -2
11(0:*  10In intervals of that much coordinate time
11                  10As often as possible
11-1  10Disable output
11-2  10Default to IO::outxdt

11

11outscalarxevery  10Scope: privatex01  INT

11Description: How often to do scalar output, overrides IO::outxevery

11 Range 10 x01 Default: -2
111:*  10Output every so many time steps
11-1:0  10No output
11-2  10Default to IO::outxevery

11

11outscalarxreductions 10Scope: privatex01  STRING

11Description: List of reductions to output in scalar form

11 Range 10 x01Default: count minimum maximum sum average norm1 norm2 normxinf icount iaverage inorm1 inorm2
11                  10A regex which matches everything

11

11outscalarxvars  10Scope: privatex01  STRING

11Description: Variables to output in scalar form

11 Range 10 x01 Default: (none)
11                  10A regex which matches everything

11

11outxcriterion  10Scope: shared from IOx01KEYWORD

11

11outxdir  10Scope: shared from IOx01STRING

11

11outxdt  10Scope: shared from IOx01REAL

11

11outxevery  10Scope: shared from IOx01INT

11

11outxfileinfo  10Scope: shared from IOx01KEYWORD

11outxgroupxseparator10Scope: shared from IOx01STRING

11

11strictxioxparameterxcheck 10Scope: shared from IOx01BOOLEAN

11

11verbose  10Scope: shared from IOx01KEYWORD

11

4 Interfaces

General Implements: ioscalar

Grid Variables

4.0.1 PRIVATE GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11lastxoutputxiteration10 lastxoutputxiteration x01  compact110
11                  10  x01 dimensions110
11                  10  x01 distribution11CONSTANT
11                  10  x01 group type11SCALAR
11                  10  x01 timelevels111
11                  10  x01variable type11INT
11 lastxoutputxtime 10 lastxoutputxtime x01  compact110
11                  10  x01 dimensions110
11                  10  x01 distribution11CONSTANT
11                  10  x01 group type11SCALAR
11                  10  x01 timelevels111
11                  10  x01variable type11REAL
11 thisxiteration 10 thisxiteration x01  compact110
11                  10  x01 dimensions110
11                  10  x01 distribution11CONSTANT
11                  10  x01 group type11SCALAR
11                  10  x01 timelevels111
11                  10  x01variable type11INT

Uses header: Timer.hh carpet.hh typecase.hh typeprops.hh

5 Schedule This section lists all the variables which are assigned storage by thorn Carpet/CarpetIOScalar. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11Always:  11x
11lastxoutputxiteration lastxoutputxtime thisxiteration11x

11x  11x

11

Scheduled Functions

CCTKxSTARTUP x carpetioscalarstartup x startup routine x

11x11After: 11ioutilxstartup
11 x 11Language:11c
11 x 11Options: 11global
11 x 11Type: 11function

11

CCTKxINITIAL x carpetioscalarinit x initialisation routine

11x11Language:11c
11 x 11Options: 11global
11 x 11Type: 11function
11 x 11Writes: 11carpetioscalar::lastxoutputxiteration
11 x 11x  11carpetioscalar::lastxoutputxtime
11 x 11x  11carpetioscalar::thisxiteration